# Quasinormal modes of Schwarzschild-de Sitter black holes in semi-open systems

### v1

@@ I Introduction
When characterizing the intrinsic oscillations of perturbed black holes, quasinormal modes (QNMs) appear as a discrete set of complex frequencies. The real part corresponds to the oscillation frequency, and the imaginary part reflects the damping rate. Serving as unique spectral fingerprints of black holes, QNMs provide a powerful probe to test the Kerr hypothesis and to investigate gravity in the strong-field regime. This field of study is referred to as black hole spectroscopy Kokkotas:1999bd; Berti:2009kk; Konoplya:2011qq; Berti:2025hly.
@@
The QNM spectra of black holes exhibit the instability Konoplya:2022pbc; Jaramillo:2020tuu; Shen:2025yiy due to the non-Hermiticity (NH) of black hole systems Ashida:2020dkc, and this phenomenon has been extensively studied. That is to say, these spectra respond with significant shifts in the complex plane to the slightest external perturbations. Initial studies of QNM spectrum instability for the black holes were provided by Nollert and Price Nollert:1996rf; Nollert:1998ys. Methods for investigating spectrum instability associated with QNMs can be classified into two primary categories. Pseudospectrum analysis trefethen2020spectra is a reliable tool to investigate the instability of the QNM spectrum Boyanov:2024fgc; Jaramillo:2020tuu; Destounis:2023ruj; Jaramillo:2021tmt. This method employs visual techniques with the goal of elucidating spectrum instability by analyzing the characteristic properties of non-self-adjoint operators in dissipative systems. Within the gravity theory, pseudospectra provide qualitative insights into spectrum instability across different spacetimes Jaramillo:2020tuu; Destounis:2021lum; Cao:2024oud; Cao:2024sot; Arean:2024afl; Garcia-Farina:2024pdd; Arean:2023ejh; Boyanov:2023qqf; Cownden:2023dam; Carballo:2025ajx; Sarkar:2023rhp; Destounis:2023nmb; Luo:2024dxl; Warnick:2024usx; Chen:2024mon; Boyanov:2022ark; Siqueira:2025lww; Besson:2024adi; dePaula:2025fqt; Cai:2025irl; Cao:2025qws. Transient dynamics associated with pseudospectra are investigated in Carballo:2024kbk; Jaramillo:2022kuv; Chen:2024mon; Carballo:2025ajx; Besson:2025ghu.
@@
However, the information provided by pseudospectra regarding spectrum stability is insufficient. This can be understood from its definition: for a given perturbation magnitude under a specific norm, the maximal spectrum shift precisely corresponds to the boundary of the pseudospectrum. While valuable, its inability to identify the precise positions of spectrum shifts greatly limits its application. Accordingly, one may consider the so-called the structured pseudospectrum trefethen2020spectra. In recent work, attention has increasingly shifted to understanding how the spectrum responds to perturbations of a given form, which are commonly added to the effective potential. Therefore, one modifies the effective potential to study the QNM spectrum instability Qian:2020cnz; Daghigh:2020jyk; Liu:2021aqh; Li:2024npg; Qian:2024iaq; Xie:2025jbr; Berti:2022xfj; Cheung:2021bol; Yang:2024vor; Courty:2023rxk; Cardoso:2024mrw; Ianniccari:2024ysv; MalatoCorrea:2025iuc; Shen:2025nsl; Wang:2025mxe; Hu:2025efp. In these references, more recently, black hole spectrum instability triggered deterministic and random metric perturbations is studied in Shen:2025nsl. For a more realistic physical model, namely Einstein-Maxwell-scalar (EMS) model, it is found that the stable and the unstable modes will coexist Wang:2025mxe. For dynamical Chern-Simons gravity, three distinctive phenomena absent in general relativity are found in terms of the QNM spectrum instability Hu:2025efp. When the small extra bump becomes negative enough, spacetime will even become unstable Mai:2025cva. The above studies are all consider within the standard QNM boundary conditions, where the ingoing boundary condition is at the event horizon and outgoing boundary condition is at infinity or the cosmological horizon.
@@
Considering that black holes in reality are not perfect, that is, they are not completely “black”, which will be replaced by the exotic compact objects (ECOs). In other words, there exists a wall with the reflectivity $\mathcal{K}(\omega)$ at $x_{0}$ in the tortoise coordinate. For these systems of ECOs, mathematically speaking, the boundary conditions around the event horizon will be modified into $\Psi\sim\mathrm{e}^{-\mathrm{i}\omega(x-x_{0})}+\mathcal{K}(\omega)\mathrm{e}^{\mathrm{i}\omega(x-x_{0})}$ at the frequency-domain. One of the resulting features of the ringdown for the ECOs is the gravitational wave (GW) echoes Mark:2017dnq; Cardoso:2019apo; Bueno:2017hyj; Cardoso:2016rao; Cao:2025qxt; Zhang:2025ygb; Yang:2025hqk. The reflectivity plays a key role in both modulating the signal and determining the amplitude of subsequent echoes. A list of models for the reflectivity has been explicitly work out such as the wormhole model Mark:2017dnq; Bueno:2017hyj, the Boltzmann model Oshita:2019sat; Wang:2019rcf, the area quantization model Deppe:2024fdo, the perfect-fluid star, gravastar, membrane paradigm models Chakraborty:2023zed. One can find a table summarizing the above models in Rosato:2025byu; Berti:2025hly. It has been universally applicable that the QNM spectra of black holes are unstable against small modifications, so even small changes in boundary conditions may lead to the QNM spectrum instability Oshita:2025ibu; Solidoro:2024yxi. Recently, ECOs version of flea and elephant models is studied in Destounis:2025dck.
@@
According to the Green’s function theory, if we can know the so-called “in” solution and “up” solution, we can calculate the (ringdown) waveforms of gravitational waves generated by any form of source in principle Andersson:1995zk; Berti:2006wq; Zhang:2013ksa; Berti:2025hly. If the “in” solution and “up” solution can be analytically solved, the technical difficulty of high-precision waveform construction will be greatly reduced. According to our research, there are Pöshl-Teller (PT) models Ferrari:1984zz, BTZ black hole Arnaudo:2025uos, Schwarzschild-de Sitter (SdS) black hole Arnaudo:2025uos; Arnaudo:2025kit, Kerr-de Sitter black hole Hatsuda:2020sbn and $\text{Kerr-AdS}_{5}$ spacetime that can be analytically solved Noda:2022zgk. Specifically, for the SdS black hole, QNMs are able to be solved by the Heun function Arnaudo:2025uos; Arnaudo:2025kit. One can refer to Chen:2025sbz for a complete treatment of QNMs of Type-D black holes, in which the Heun function is also used. In addition, for the Schwarzschild black hole, the Coulomb wave functions are used to construct the Green function Leaver:1986gd; Berti:2025hly. More recently, in the Grumiller spacetime Mi:2025fbt and in the rotating Kalb-Ramond BTZ black hole Xia:2025hwt, QNMs can be solved analytically.
@@
In this study, we focus on the perturbations of the SdS black hole. The SdS spacetime models a non-spinning, uncharged black hole in an asymptotically de Sitter universe, being a black hole solution with a positive cosmological constant $\Lambda>0$. Its astrophysical importance arises from cosmological evidence for a positive cosmological constant, meaning real astrophysical black holes reside in spacetimes where this constant, though small, modifies the large-scale asymptotic geometry. The QNM spectra of SdS black hole have been computed in Zhou:2025xta; Cardoso:2003sw; Konoplya:2004uk; Jansen:2017oag; Cardoso:2017soq; Konoplya:2022xid by pseudo-spectral method, Leaver method or near-extreme approximation method, respectively. One of our studies in this work is to show how the spectra of the SdS black hole are affected by reflectivity $\mathcal{K}(\omega)$ being near the event horizon.
@@
While frequency-domain analyses reveal that small disturbances can cause substantial spectral changes, as demonstrated by studies of effective potential modifications, pseudospectrum approaches and boundary condition modifications, the ringdown waveform remains stable Wu:2025sbq; Daghigh:2020jyk; Berti:2022xfj; Yang:2024vor; Oshita:2025ibu; Kyutoku:2022gbr. The stability of waveform and the stability of greybody factor (GF) are inseparable Kyutoku:2022gbr. In other words, the black hole greybody factors are more robust observables Oshita:2023cjz; Rosato:2024arw; Oshita:2024fzf; Ianniccari:2024ysv; Wu:2024ldo; Kyutoku:2022gbr; Konoplya:2025ixm; Wu:2025sbq; Xie:2025jbr unlike the QNM spectra. Although QNMs and GFs are both characteristic quantities reflecting the geometry of spacetime, they are defined under different boundary conditions. Their correspondence was first established for spherically symmetric black holes in Konoplya:2024lir. The correspondence between QNMs and GFs has been extended to various black holes Dubinsky:2025nxv; Konoplya:2024vuj; Bolokhov:2024otn; Skvortsova:2024msa; Malik:2024cgb. For ultracompact horizonless objects, Refs. Rosato:2025byu; Rosato:2025lxb give detail discussion on the GFs. Therefore, motivated by them, we also investigate the GFs on the SdS black hole within the existence of the reflectivity. Here, we focus on two cases. One is that $\mathcal{K}(\omega)$ is a real constant independent of the frequency, while the other is that $\mathcal{K}(\omega)$ is the Boltzmann reflectivity Oshita:2019sat; Wang:2019rcf.
@@
Black holes are non-Hermiticity (NH) systems. Such NH systems exist in many fields of physics Ashida:2020dkc. A hallmark of NH systems is the existence of exceptional point (EP) Heiss:1999alz; Heiss:2012dx; Ryu:2025gnn; Zhong:2018hmw; Dietz:2010bvm, defined as a spectral singularity where both eigenvalues and their corresponding eigenvectors coalesce. For more mathematical aspects of EPs, we can refer to monograph kato1976perturbation. For the gravity aspects, EPs also appear in the QNM framework of black hole perturbation systems. For massive scalar perturbation of a Kerr black hole, two QNM spectra can become degenerate Cavalcante:2024swt; Cavalcante:2024kmy; Cavalcante:2025abr, where the EP is $(M\mu)_{c}\simeq 0.3704981$ and $(a/M)_{c}\simeq 0.9994660$. Such EP is also obtained in Chen:2025sbz. The EPs can also be found when a Gaussian bump is added to the Regge-Wheeler effective potential Yang:2025dbn, where the EP is $\epsilon\simeq 0.005$, and $d\simeq 15.698$. Near such EPs, the conventional decomposition of the ringdown waveform in to a superposition of QNMs may become invalid Yang:2025dbn; PanossoMacedo:2025xnf. A phenomenon related to the concept of exceptional points is called mode repulsion or avoided crossing, which have been widely studied in Dias:2021yju; Motohashi:2024fwt; Oshita:2025ibu; Lo:2025njp; Takahashi:2025uwo. When the dimension of parameters increases, EPs will form exception lines (ELs) which widely arises in non-Hermitian systems Shen:2018cjc; Yang:2018dzr; Wu:2024wlf. For the gravity aspect, in Cao:2025afs, EL is found, and the topological structures (such Berry phase and vorticity) of QNMs associated with EL are also considered. Although we know that the spectrum is unstable, the spectra near exceptional points will be even more unstable, which is studied through the pseudospectra at exceptional points in Cao:2025afs. In this study, we change the boundary conditions to find exceptional points in the SdS black hole. As far as we know, it is the first work of EPs associated with boundary condition changing in the framework of gravity aspect. More specifically, the reflectivity $\mathcal{K}$ can be prompted into a complex parameter, $\mathcal{K}\in\mathbb{C}$, thus provides two variable real parameters. The existence of imaginary parts in reflectivity means that when waves are reflected at the boundary around the event horizon, not only does their amplitude change, but their phase varies as well. This phase change thereby systematically shifting its QNM spectra.
@@
The work is organized as follows. In Sec. II, we solve the perturbation equation of SdS black hole by the Heun function, and give the condition of the QNM spectra as the reflectivity exists. In Sec. III, the migration of the QNM spectra with respect to the reflectivity in the near-extremal and near-Schwarzschild cases is studied, in which the reflectivity $\mathcal{K}$ restricts to be real and $\mathcal{K}\in[0,1]$. In Sec. IV, the influence of reflectivity on the greybody factor is explored. In Sec. V, considering that reflectivity is a complex constant, $\mathcal{K}\in\mathbb{C}$, we find out the exception point. Sec. VI is the conclusions and discussion. In Appendix A, the process of using the Heun functions to solve QNMs in the SdS black hole is shown. Four expressions of amplitude functions, $A_{\text{in}}(\omega)$, $A_{\text{out}}(\omega)$, $B_{\text{in}}(\omega)$ and $B_{\text{out}}(\omega)$, are given in Appendix B.

### v2

@@ I Introduction
When characterizing the intrinsic oscillations of perturbed black holes, quasinormal modes (QNMs) appear as a discrete set of complex frequencies. The real part corresponds to the oscillation frequency, and the imaginary part reflects the damping rate. Serving as unique spectral fingerprints of black holes, QNMs provide a powerful probe to test the Kerr hypothesis and to investigate gravity in the strong-field regime. This field of study is referred to as black hole spectroscopy Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011); Berti and others (2025).
@@
The QNM spectra of black holes exhibit the instability Konoplya and Zhidenko (2024b); Jaramillo et al. (2021); Shen et al. (2025a) due to the non-Hermitian (NH) of black hole systems Ashida et al. (2021), and this phenomenon has been extensively studied. That is to say, these spectra respond with significant shifts in the complex plane to the slightest external perturbations. Initial studies of QNM spectrum instability for the black holes were provided by Nollert and Price Nollert (1996); Nollert and Price (1999). Methods for investigating spectrum instability associated with QNMs can be classified into two primary categories. Pseudospectrum analysis Trefethen and Embree (2005) is a reliable tool to investigate the instability of the QNM spectrum Boyanov (2024); Jaramillo et al. (2021); Destounis and Duque (2023); Jaramillo et al. (2022). This method employs visual techniques with the goal of elucidating spectrum instability by analyzing the characteristic properties of non-self-adjoint operators in dissipative systems. Within the gravity theory, pseudospectra provide qualitative insights into spectrum instability across different spacetimes Jaramillo et al. (2021); Destounis et al. (2021); Cao et al. (2024, 2025c); Arean et al. (2024); Garcia-Fariña et al. (2025); Areán et al. (2023); Boyanov et al. (2024); Cownden et al. (2024); Carballo et al. (2025); Sarkar et al. (2023); Destounis et al. (2024); Luo (2024); Warnick (2024); Chen et al. (2024); Boyanov et al. (2023); Siqueira et al. (2025); Besson and Jaramillo (2025); de Paula et al. (2025); Cai et al. (2025); Cao et al. (2025b). Transient dynamics associated with pseudospectra are investigated in Carballo and Withers (2024); Jaramillo (2022); Chen et al. (2024); Carballo et al. (2025); Besson et al. (2025).
@@
However, the information provided by pseudospectra regarding spectrum stability is insufficient. This can be understood from its definition: for a given perturbation magnitude under a specific norm, the maximal spectrum shift precisely corresponds to the boundary of the pseudospectrum. While valuable, its inability to identify the precise positions of spectrum shifts greatly limits its application. Accordingly, one may consider the so-called structured pseudospectrum Trefethen and Embree (2005). In recent work, attention has increasingly shifted to understanding how the spectrum responds to perturbations of a given form, which are commonly added to the effective potential. Therefore, one modifies the effective potential to study the QNM spectrum instability Qian et al. (2021); Daghigh et al. (2020); Liu et al. (2021); Li et al. (2024); Qian et al. (2025); Xie et al. (2025); Berti et al. (2022); Cheung et al. (2022); Yang et al. (2024); Courty et al. (2023); Cardoso et al. (2024); Ianniccari et al. (2024); Malato Corrêa et al. (2025); Shen et al. (2025b); Wang and Wu (2025); Hu et al. (2025). In these references, more recently, black hole spectrum instability triggered deterministic and random metric perturbations is studied in Shen et al. (2025b). For a more realistic physical model, namely Einstein-Maxwell-scalar (EMS) model, it is found that the stable and the unstable modes will coexist Wang and Wu (2025). For dynamical Chern-Simons gravity, three distinctive phenomena absent in general relativity are found in terms of the QNM spectrum instability Hu et al. (2025). When the small extra bump becomes negative enough, spacetime will even become unstable Mai and Yang (2025). The above studies are all considered within the standard QNM boundary conditions, where the ingoing boundary condition is at the event horizon and outgoing boundary condition is at infinity or the cosmological horizon.
@@
Considering that black holes in reality are not perfect, that is, they are not completely “black”, which will be replaced by the exotic compact objects (ECOs). In other words, there exists a wall with the reflectivity $\mathcal{K}(\omega)$ at $x_{0}$ in the tortoise coordinate. For these systems of ECOs, mathematically speaking, the boundary conditions around the event horizon will be modified into $\Psi\sim\mathrm{e}^{-\mathrm{i}\omega(x-x_{0})}+\mathcal{K}(\omega)\mathrm{e}^{\mathrm{i}\omega(x-x_{0})}$ at the frequency-domain. One of the resulting features of the ringdown for the ECOs is the gravitational wave (GW) echoes Mark et al. (2017); Cardoso et al. (2019); Bueno et al. (2018); Cardoso et al. (2016); Cao et al. (2025d); Zhang et al. (2025); Yang et al. (2025b). The reflectivity plays a key role in both modulating the signal and determining the amplitude of subsequent echoes. A list of models for the reflectivity has been explicitly work out such as the wormhole model Mark et al. (2017); Bueno et al. (2018), the Boltzmann model Oshita et al. (2020); Wang et al. (2020), the area quantization model Deppe et al. (2025), the perfect-fluid star, gravastar, membrane paradigm models Chakraborty et al. (2024). One can find a table summarizing the above models in Rosato et al. (2025b); Berti and others (2025). It has been universally applicable that the QNM spectra of black holes are unstable against small modifications, so even small changes in boundary conditions may lead to the QNM spectrum instability Oshita et al. (2025); Solidoro et al. (2025). Recently, ECOs version of flea and elephant models has been studied in Destounis et al. (2025).
@@
According to the Green’s function theory, if we can know the so-called “in” solution and “up” solution, we can calculate the (ringdown) waveforms of gravitational waves generated by any form of source in principle Andersson (1995); Berti and Cardoso (2006); Zhang et al. ; Berti and others (2025). If the “in” solution and “up” solution can be analytically solved, the technical difficulty of high-precision waveform construction will be greatly reduced. According to our research, there are Pöshl-Teller (PT) models Ferrari and Mashhoon (1984), BTZ black hole Arnaudo et al. (2025), Schwarzschild-de Sitter (SdS) black hole Arnaudo et al. (2025); Arnaudo and Withers (2025), Kerr-de Sitter black hole Hatsuda (2020) and $\text{Kerr-AdS}_{5}$ spacetime that can be analytically solved Noda and Motohashi (2022). Specifically, for the SdS black hole, QNMs are able to be solved by the Heun function Arnaudo et al. (2025); Arnaudo and Withers (2025). One can refer to Chen et al. (2025a) for a complete treatment of QNMs of Type-D black holes, in which the Heun function is also used. In addition, for the Schwarzschild black hole, the Coulomb wave functions are used to construct the Green function Leaver (1986); Berti and others (2025). More recently, in the Grumiller spacetime Mi and Li (2025) and in the rotating Kalb-Ramond BTZ black hole Xia et al. (2025), QNMs can be solved analytically.
@@
In this study, we focus on the perturbations of the SdS black hole. The SdS spacetime models a non-spinning, uncharged black hole in an asymptotically de Sitter universe, being a black hole solution with a positive cosmological constant $\Lambda>0$. Its astrophysical importance arises from cosmological evidence for a positive cosmological constant, meaning real astrophysical black holes reside in spacetimes where this constant, though small, modifies the large-scale asymptotic geometry. The QNM spectra of SdS black hole have been computed in Zhou and Panosso Macedo (2025); Cardoso and Lemos (2003); Konoplya and Zhidenko (2004); Jansen (2017); Cardoso et al. (2018); Konoplya and Zhidenko (2022) by pseudo-spectral method, Leaver method or near-extreme approximation method, respectively. One of our studies in this work is to show how the spectra of the SdS black hole are affected by reflectivity $\mathcal{K}(\omega)$ being near the event horizon.
@@
While frequency-domain analyses reveal that small disturbances can cause substantial spectral changes, as demonstrated by studies of effective potential modifications, pseudospectrum approaches and boundary condition modifications, the ringdown waveform remains stable Wu et al. (2025b); Daghigh et al. (2020); Berti et al. (2022); Yang et al. (2024); Oshita et al. (2025); Kyutoku et al. (2023). The stability of waveform and the stability of greybody factor (GF) are inseparable Kyutoku et al. (2023). In other words, the black hole greybody factors are more robust observables Oshita (2024); Rosato et al. (2024); Oshita et al. (2024); Ianniccari et al. (2024); Wu et al. (2025a); Kyutoku et al. (2023); Konoplya and Pappas (2025); Wu et al. (2025b); Xie et al. (2025) unlike the QNM spectra. Although QNMs and GFs are both characteristic quantities reflecting the geometry of spacetime, they are defined under different boundary conditions. Their correspondence was first established for spherically symmetric black holes in Konoplya and Zhidenko (2024a). The correspondence between QNMs and GFs has been extended to various black holes Dubinsky (2026); Konoplya and Zhidenko (2025); Bolokhov and Skvortsova (2025); Skvortsova (2025); Malik (2025). For ultracompact horizonless objects, Refs. Rosato et al. (2025b, a) give detailed discussion on the GFs. Therefore, motivated by them, we also investigate the GFs on the SdS black hole within the existence of the reflectivity. Here, we focus on two cases. One is that $\mathcal{K}(\omega)$ is a real constant independent of the frequency, while the other is that $\mathcal{K}(\omega)$ is the Boltzmann reflectivity Oshita et al. (2020); Wang et al. (2020).
@@
Black holes are non-Hermitian (NH) systems. Such NH systems exist in many fields of physics Ashida et al. (2021). A hallmark of NH systems is the existence of exceptional point (EP) Heiss (1999, 2012); Ryu et al. (2025); Zhong et al. (2018); Dietz et al. (2011), defined as a spectral singularity where both eigenvalues and their corresponding eigenvectors coalesce. For more mathematical aspects of EPs, we can refer to monograph Kato (1976). For the gravity aspects, EPs also appear in the QNM framework of black hole perturbation systems. For massive scalar perturbation of a Kerr black hole, two QNM spectra can become degenerate Cavalcante et al. (2024a, b, 2025), where the EP is $(M\mu)_{c}\simeq 0.3704981$ and $(a/M)_{c}\simeq 0.9994660$. Such EP is also obtained in Chen et al. (2025a). The EPs can also be found when a Gaussian bump is added to the Regge-Wheeler effective potential Yang et al. (2025a), where the EP is $\epsilon\simeq 0.005$, and $d\simeq 15.698$. Near such EPs, the conventional decomposition of the ringdown waveform into a superposition of QNMs may become invalid Yang et al. (2025a); Panosso Macedo et al. (2025). A phenomenon related to the concept of exceptional points is called mode repulsion or avoided crossing, which have been widely studied in Dias et al. (2022); Motohashi (2025); Oshita et al. (2025); Lo et al. (2025); Takahashi et al. (2025). When the dimension of parameters increases, EPs will form exceptional lines (ELs) which widely arises in non-Hermitian systems Shen et al. (2018); Yang and Hu (2019); Wu et al. (2024). For the gravity aspect, in Cao et al. (2025a), EL is found, and the topological structures (such Berry phase and vorticity) of QNMs associated with EL are also considered. Although we know that the spectrum is unstable, the spectra near exceptional points will be even more unstable, which is studied through the pseudospectra at exceptional points in Cao et al. (2025a).
In this study, we change the boundary conditions to find exceptional points in the SdS black hole, namely by introducing a complex reflectivity $\mathcal{K}$. In principle, EPs could also be searched for in other two-parameter spaces, e.g. by varying $(\mathcal{K}\in\mathbb{R},\,r_{c})$, where $r_{c}$ corresponds to the cosmological constant (refer to Eqs. (2)). However, we do not adopt it because changing $r_{c}$ modifies the background spacetime (i.e. the cosmological constant), whereas our goal is to isolate EPs as an intrinsic effect induced by the semi-open boundary condition itself. As far as we know, it is the first work of EPs associated with boundary condition changing in the framework of gravity aspect. More specifically, the reflectivity $\mathcal{K}$ can be prompted into a complex parameter, $\mathcal{K}\in\mathbb{C}$, thus provides two variable real parameters. The existence of imaginary parts in reflectivity means that when waves are reflected at the boundary around the event horizon, not only does their amplitude change, but their phase varies as well. This phase change thereby systematically shifting its QNM spectra.
@@
The work is organized as follows. In Sec. II, we solve the perturbation equation of SdS black hole by the Heun function, and give the condition of the QNM spectra as the reflectivity exists. In Sec. III, the migration of the QNM spectra with respect to the reflectivity in the near-extremal and near-Schwarzschild cases is studied, in which the reflectivity $\mathcal{K}$ restricts to be real and $\mathcal{K}\in[0,1]$. In Sec. IV, the influence of reflectivity on the greybody factor is explored. In Sec. V, considering that reflectivity is a complex constant, $\mathcal{K}\in\mathbb{C}$, we find out the exception point. Sec. VI is the conclusions and discussion. In Appendix A, the process of using the Heun functions to solve QNMs in the SdS black hole is shown. Four expressions of amplitude functions, $A_{\text{in}}(\omega)$, $A_{\text{out}}(\omega)$, $B_{\text{in}}(\omega)$ and $B_{\text{out}}(\omega)$, are given in Appendix B. A concrete result of QNMs for $s=0$ is shown in Appendix C.